Use the convolution theorem to find the inverse Laplace transform of the given function. 2 $3 (s2+4) 2 >(t)= (s2+4) s3
Answer #1

The convolutim of flt) and gets is fing) (A) (4-0) gires dre. L { Cfxg) legy = f(s) GCS). = L{fix)} {gles} f(s) :{fiksy G(S)(fagle) = (t-ug? Sinzo dr. 2 (t in 20-210. in 20 tu? Sinzu) dhe ta Cosau Now, s t2 dinzo de 2 fate sinon Sztv.om 28 3-240 692